A bookcase with three shelves is to be built by a woodworking student. If the height of the bookcase is to be 2 feet longer than the length of a shelf and the total amount of to be used is 32 feet, find the dimensions of the bookcase.
I have tried using 32=4L+2h+2

Answer by macston(5194)

L=length of shelf; H=height=L+2
.
2 pieces (sides)- L+2ft long each
3 pieces (shelves) - L long each
.
32 feet=2(sides)+3(shelves)
32 feet=2(L+2ft)+3(L)
32ft=2L+4ft+3L
32ft=5L+4ft
28ft=5L
5.6ft=L
H=L+2ft=5.6ft+2ft=7.6ft
.
ANSWER: The bookcase is 7.6 feet high by 5.6 feet wide.
.